OpenStudy (vortish):

-3/8x + 1/2 =3/8 solve and check the equation

OpenStudy (anonymous):

you have some options... you could start by multiplying everything through by 8... that would get rid of the 8's in the denominators.

OpenStudy (ny,ny):

or you could subtract by 1/2 on both sides...

OpenStudy (anonymous):

that's also a good first step.

OpenStudy (vortish):

ok so if i multiplied every thing by 8 we would get 24/8 8/8 or 1 and 24/8 correct

OpenStudy (anonymous):

I agree, subtract by 0.5 on both sides, leaves you with -(1/8). Multiply both sides by 8x and that should do it.

hero (hero):

\[-\frac{3x}{8} + \frac{1}{2} =\frac{3}{8}\] \[ \frac{1}{2} =\frac{3}{8} +\frac{3x}{8}\] \[ \frac{1}{2} =\frac{3+3x}{8}\] Cross Multiply: 8 = 2(3 + 3x) 8 = 6 + 6x 8 - 6 = 6x 2 = 6x 2/6 = x 1/3 = x

OpenStudy (vortish):

im not looking for the answer just the steps to come up with the answer 3/8= 3/8 - 1/2

hero (hero):

I provided steps

OpenStudy (vortish):

hero that makes sense thanks for all the help all of you guys are great wish I understood this stuff better